OpenStudy (stamp):

Use the following function \[g(x)=x^3-(-3x^2)-9x-7\] Determine the absolute maxima and the absolute minima for g(x) on the interval [1, 5].

OpenStudy (stamp):

\[g(1)=1+1-9-7=2-16=-14\]\[g(5)=125+75-45-7=200-52=148\] Absolute maxima and minima on the interval [1, 5] is maxima (1, -14) and minima (5, 148)

OpenStudy (stamp):

@Loser66 logic check me, I am hoping I interpreted the question correctly